OpenStudy (anonymous):

help please 50/x / 100/x^2+5x

OpenStudy (anonymous):

\[\frac{50}{x}\times \frac{x^2+5x}{100}\] is a first step then factor and cancel

OpenStudy (anonymous):

\[\frac{50x(x+5)}{100x}=\frac{x+5}{2}\]

OpenStudy (anonymous):

that's the answer??